OpenStudy (anonymous):

solve 3 over y = y over 5y-12

OpenStudy (anonymous):

3/y = y /5y-12 y^2 = 15y - 36 y^2 - 15y + 36 = 0 (y-3)(y-12) = 0 solution 3 , 12